A 3kg object travels in a uniform circle at a constant speed of 12m/s. The centripetal force acting on the object is 110 Newtons.


Calculate the weight of the object


1
Expert's answer
2022-01-30T13:52:20-0500

Given:

"m=3\: \rm kg"

"v=12\: \rm m/s"

"F_c=110\:\rm N"


The weight of an object

"W=m(a+g)=F_c+mg\\
\rm =110\: N+3\: kg*9.8\: m/s^2=139.4\: N"